OpenStudy (anonymous):

Let F(x)=integral of sin(t^2) dt from 0 to x for 0<=x<=3. On what interval or intervals is F increasing? Justify your answer.

OpenStudy (anonymous):

I know that F'(x)=sin(x^2).

OpenStudy (anonymous):

But what to do next?

OpenStudy (anonymous):

set it equal to 0 and find solve for x. Then use number line test

OpenStudy (anonymous):

sin(x^2)=0 but how do I solve for x though?

OpenStudy (anonymous):

when is sine 0?

OpenStudy (anonymous):

0, pi, 2pi.

OpenStudy (anonymous):

techically it's 0, pi, 2pi, 3pi , .... k pi, and k is an integer. But it turns out that only 0, pi and 2pi work in this case. so x^2 = 0, pi, 2pi then x = 0, sqrt(pi), sqrt(2pi)

OpenStudy (anonymous):

if your domain was, say, [0,10], then you might have to include sqrt(3pi), sqrt(4pi), ... and whichever one that is in that range.

OpenStudy (anonymous):

So what's the answer for this problem?

OpenStudy (anonymous):

so you have to check is F'(x) positive in (0,sqrt(pi)) ? is F'(x) positive in (sqrt(pi), sqrt(2pi)? is F'(x) positive in (sqrt(2pi), 3)? pick a number that is in each of those interval and plug it in F'(x) and see if you get a positive number. If you do, F(x) is increasing during that interval

OpenStudy (anonymous):

Thanks.